OpenStudy (anonymous):

lim x>0 (cube root(1+2x) -1)/x

OpenStudy (anonymous):

\[\frac{ \sqrt[3]{1+2x} -1}{ x }\]

OpenStudy (phi):

do you know l'hospital ?

OpenStudy (anonymous):

yes, please if possible it's not in our module so the answer wont be acceptable

OpenStudy (turingtest):

make it into a difference of cubes

OpenStudy (turingtest):

multiply on top and below the right factors to give the formula (a-b)(a^2+ab+b^2) in your case\[a=\sqrt[3]{1+2x},~~b=1\]

OpenStudy (anonymous):

thanx @TuringTest . let me see

OpenStudy (anonymous):

am stuck here @TuringTest , whi ch ones are the right factors

OpenStudy (turingtest):

if\[a=\sqrt[3]{1+2x},~~b=1\]then\[a^2+ab+b^2=(1+2x)^{2/3}+(1+2x)^{1/3}+1\]multiply that above and below, and usethe identity \[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

i have done that, question is,now, what is next? what should i do?

OpenStudy (turingtest):

what are you left with?

OpenStudy (turingtest):

Show what you have so far please, so I can help you. If you did what I told you the cancellations should be obvious and you can just take the limit.

OpenStudy (anonymous):

i am sending a pic

OpenStudy (anonymous):

hope you can see it @TuringTest

OpenStudy (turingtest):

that's correct, now use the identity \[a^3-b^3=(a-b)(a^2+ab+b^2)\]on the top

OpenStudy (raden):

this is just an alternative, hope can help you too let u = 1+2x -----> x = (u-1)/2 if x->0 then u -> 1 so, the limit becomes lim u->1 (u^1/3 - 1)/((u-1)/2) = lim u->1 2*(u^1/3 - 1)/(u-1) factor out the denominator, get = lim u->1 2*(u^1/3-1)/[(u^1/3-1)(u^2/3+u^1/3+1)] now you can cancel out the similar terms, get = lim u->1 2/(u^2/3+u^1/3+1) finally, just subtitute u = 1 now, get your answer!

OpenStudy (anonymous):

thanx @RadEn - i will definatly pursue it.. I am equally interested in the method at hand. so which one becomes my a and b on the numerator @TuringTest . should i fatorise or expand? or is \[(1 +2x)^{\frac{ 1 }{ 3}} -1)\] =a and the other parenthesis b ?

OpenStudy (anonymous):

on the numerator

OpenStudy (turingtest):

no expanding, just run the identity as I wrote it from right to left\[{(1+2x)^{1/3}[(1+2x)^{2/3})+(1+2x)^{1/3}+1]\over x[(1+2x)^{2/3})+(1+2x)^{1/3}+1]}\]noticing that the numerator is of the form\[(a-b)(a^2+ab+b^2)\] with \(a=(1+2x)^{1/3}\) and \(b=1\) we can rewrite it in the form \(a^3-b^3\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{(1+2x)^{\frac{ 2 }{ 3}}-1) }{ x \left[ (1+2x)^{\frac{ 2 }{ 3 }}+(1+2x)^{\frac{ 1 }{ 3 } }+1 \right] }\] yeh am not gettin it

OpenStudy (turingtest):

how did you get that numerator? what is a? what is a^3 what is b? what is b^3?

OpenStudy (anonymous):

but this is big stuff so i want to know it

OpenStudy (anonymous):

honestly, i just took the b you gave to me and squared it, same with the a

OpenStudy (anonymous):

i know the difference of ubes, but i am failing to figure it out in this one

OpenStudy (turingtest):

why square it? we are using difference of *cubes*. difference of squares looks totally different, you ought to memorize those

OpenStudy (turingtest):

\[a^3-b^3=(a-b)(a^2+ab+b^2)\] what is a? what is a^3?

OpenStudy (anonymous):

oh wait!

OpenStudy (anonymous):

@TuringTest - u are a genius man. I have realised that the numerator will be 1 +2x-1 =2x, the x cancels with the one on the denominator . this leaves us with 2 on top, and then we plug in zero to get 1 + 1+ 1 , so 2/3 is the answer thanx!

OpenStudy (turingtest):

welcome! These kinds of problems often have people going straight to l'hospital, bit l'hospital is actually invalid in this case anyway. either way, be sure to try out all the little algebra tricks that seemed useless (diff square, diff cubes, complete square, etc). That's usually the ticket ;)

OpenStudy (anonymous):

thanx a million times. have just learnt something SERIOUSLY NEW today!

OpenStudy (turingtest):

that's better than a million medals :)

OpenStudy (anonymous):

@RadEn